OpenStudy (anonymous):

circle A has equation (x+5)^2+y^2=169. The diameter of circle B is one fourth as long as the diameter of circle A. What is the radius of circle B?

OpenStudy (campbell_st):

the general form for a circle is \[(x -h)^2 + (y - k)^2 = r^2\] in your question the centre is (-5, 0) so h = -5 and k = 0 and r^2 = 169 to find the radius, take the square root. to find the radius of circle B divide your radius by 4

OpenStudy (anonymous):

divide 169 by 4 or do I divide the square root of 169's square root by 4? @campbell_st

OpenStudy (campbell_st):

no divide the square root of 169 by 4 the square root of 169 is the radius of circle A

OpenStudy (anonymous):

okay thank you so much! @campbell_st